Leader: Bryan Moffit
Target: Brandon Craver
Beam Energy = 0.362 GeV
Summary
Nice data until an event in the injector left us beam-less.
Significant Events
- 08:00 Began shift in the middle of
runs 6193/26181.
- 08:10 Bob Michaels calls to see if
situation with deadtime issue has changed. May come in later today.
- 08:28 Move to carbon target.
- 09:55 Status check from Brad.
- 10:13 Move to B4C target.
- 11:09 Move to Li target.
- 12:11 Alarm on a couple channels in the
beamline HV crate. Card #1: Ch 0 and 1 tripped off. I turned them
back on.
- 12:15 Moved left spectrometer to 51.0°
- 12:30 MCC says the gun tripped. Looking
into solving this problem.
- 13:30 Controlled access.
- 13:55 ARM and Alexander return from the
Hall.
- 15:35 MCC calls with an update on the Gun
troubles. Could create a significant delay in beam delivery (few
hours to a couple of days).
Data Run Summary
Left Arm θ = 46°, Right Arm θ = 24°
- 08:00 Run
6193/26181 - Ta Target. 2.0x2.0 raster. 15 µA.
- 08:32 Run
6194/26182 - C Target. 2.0x2.0 raster. 5 µA.
- 09:22 Run
6195/26183 - C Target. 2.0x2.0 raster. 5 µA.
- 10:15 Run
6196/26184 - B4C Target. 2.0x2.0 raster. 5
µA. Wrong Prescales.
- 10:21 Run
6197/26185 - B4C Target. 2.0x2.0 raster. 5
µA. Wrong Prescales.
- 10:32 Run
6198/26186 - B4C Target. 2.0x2.0 raster. 5
µA.
- 10:51 Run
6199/26187 - B4C Target. 2.0x2.0 raster. 5
µA.
- 11:18 Run
6200/26188 - Li Target. 2.0x2.0 raster. 5 µA. Wrong
prescales.
- 11:21 Run
6201/26189 - Li Target. 2.0x2.0 raster. 5 µA.
- 11:49 Run
6202/26190 - Li Target. 2.0x2.0 raster. 5 µA.
Several CODA runs to diagnose DAQ deadtime issues
|